### v1

# A Picture Is Worth a Graph: Blueprint Debate on Graph for Multimodal Reasoning

Abstract.
This paper presents a pilot study aimed at introducing multi-agent debate into multimodal reasoning. The study addresses two key challenges: the trivialization of opinions resulting from excessive summarization and the diversion of focus caused by distractor concepts introduced from images. These challenges stem from the inductive (bottom-up) nature of existing debating schemes. To address the issue, we propose a deductive (top-down) debating approach called Blueprint Debate on Graphs (BDoG). In BDoG, debates are confined to a blueprint graph to prevent opinion trivialization through world-level summarization. Moreover, by storing evidence in branches within the graph, BDoG mitigates distractions caused by frequent but irrelevant concepts. Extensive experiments validate BDoG, achieving state-of-the-art results in Science QA and MMBench with significant improvements over previous methods.
1. Introduction
Multimodal reasoning depends on two key aspects: the creation of a unified representation of semantics from different modalities and the integration of these diverse semantics while ensuring logical consistency. While the advancement in large language models (LLMs) has made it possible to represent the semantics in natural languages (Achiam et al., 2023), the integration of diverse semantics remains a challenging issue, even in exclusive NLP tasks. One approach to tackle this challenge is multi-agent debate (MAD), where multiple LLMs act as agents, each contributing their own perspectives on the target topic and reaching a consensus through debates (Liang et al., 2023; Chan et al., 2023). This scheme could be adopted by incorporating a specific LLM for each modality as an agent.
Despite being relatively unexplored in the multimodal domain, MAD encounters numerous challenges in a broader context. MAD may suffer from the trivialization of opinions, which is a result of the summarization step performed at the conclusion of each debating round. The objective of this step is to seek agreement among the participating agents regarding their opinions. Consequently, this process can lead to the debate’s focus being directed towards a general concept, serving as an adaptation to accommodate the diverse range of semantics. One example can be observed in the reasoning of the Multimodal Language Model (MLLM) depicted in Figure 1, where the image modality presents a diverse range of semantics, including bear sedge, earthworm, collared lemming, and others. As a consequence, this can result in the context and summary being trivialized, shifting the emphasis from lichen to a more generalized concept of the tundra ecosystem, wherein both bilberry and mushroom exhibit a high degree of correlation. Similar issue exists when MAD is employed, where the summarizer concludes the diverse semantics into general words like ecosystem and food web, making the conclusion less specific. In addition, MAD may encounter the issue of focus diversion, which occurs when Chain-of-Thoughts (CoT) is utilized and new concepts introduced are highly correlated with a particular concepts (e.g., mathematical model (Cobbe et al., 2021)), leading to an increased weighting of that concept within the context.
We argue that these challenges arise due to the inductive nature of existing debating schemes, wherein agent opinions are gathered from disparate concepts at word-level and consensus is achieved through bottom-up summarization. This approach may be effective in confined NLP tasks (He et al., 2020; Hendrycks et al., 2020), where the topic is often limited to a small number of concepts and the application of CoT remains constrained. However, in a multimodal scenario, certain modalities (e.g., images) are information-rich and have a higher likelihood of introducing distracting concepts (Lu et al., 2022). Consequently, it increases the semantic divergence within the context and the likelihood of trivialization. The semantic divergence increases further when the impacts of those concepts are amplified through CoT, particularly when the newly introduced concepts exhibit biases towards certain concepts, resulting in focus diversion.
To address this issue, we propose an deductive reasoning scheme called Blueprint Debate on Graph (BDoG, pronounced bee-dog). BDoG is inspired by the blueprint debate that has been widely employed in real-world debates, which distinguishes itself from other debates by its concentration on evaluating and refining a proposal (e.g., blueprint) to address specific challenges or issues. BDoG begins by aggregating concepts from modalities and incorporating with their relationships into a initial graph. This graph serves as a blueprint that confines the scope of the discussion rather than having it open to irrelevant semantics as in existing schemes. More importantly, BDoG conducts the debate in a top-down manner by marking down conclusions on the graph. This prevents trivialization as specific concepts are preserved rather than merged into general ones. This can be found from the example shown in Figure 1, where the scope is limited to the tundra ecosystem while specific concepts like mushroom and lichen are retained. Furthermore, the graph provide a compact and high-level guidance for the discussion process. The newly introduced concepts are incorporated into relevant branches instead of remaining as a word-level thoughts within the context. This reduces the likelihood of focus diversion since, in DBoG, the competition of semantics occurs at the branch level rather than the word level. This can be seen from Figure 1, where the most relevant branches related to the soil and caribou standout from the competition, eliminating the irrelevant semantcis effectively. In addition to the advantages of scope-confined guidance and branch-level competition, BDoG also increases explainability, allowing for the tracking of discussion progress. This is evident in Figure 1, in which the progress of debating is more understandable than word-level reasoning (i.e., MLLM-CoT or MAD).
2. Related Works
2.1. Multimodal Reasoning
Multimodal reasoning is a crucial component in the development of advanced artificial intelligence (AI) systems that aim to replicate human-like intelligence (Lu et al., 2022). This type of reasoning enables AI systems to process and analyze information from various sources and forms, such as text, images, audio, and video, in a integrated and coordinated manner (Nam et al., 2017; Cadene et al., 2019). The latest advancements in multimodal large language models, such as BLIP2 (Li et al., 2023a), KOSMOS (Huang et al., 2024) and LLaVA (Liu et al., 2024) have demonstrated significant progress in complex reasoning, as these models (Zhang et al., 2023b) now have the capability to generate step-by-step rationales prior to producing the final answer, following a chain-of-thought (CoT) manner. Zheng et al. (Zheng et al., 2023b) propose a duty-distinct prompting method wherein questions are decomposed into sub-questions to enable deep-layer reasoning. SCITUNE (Horawalavithana et al., 2023) and T-SciQ (Wang et al., 2023b) aim to teach large language models to answer science questions via the generation of mixed rationales derived from both large pretrained models and human annotators. Chameleon (Lu et al., 2024) accomplishes complex multimodal reasoning tasks by integrating various external tools (e.g., large language models, off-the-shelf vision models, and web search engines). Nevertheless, existing methods exhibit limitations as they heavily rely on either few-shot learning or supervision to “guide” the reasoning process. In order to overcome this dependency, we propose the incorporation of debating features into our method. This enables the agents to engage in an adversarial discussion, allowing them to “figure out” the correct direction autonomously. As a result, our approach makes zero-shot learning feasible by reducing the reliance on external guidance or supervision.
2.2. Multi-agent Debate
To mitigate the susceptible error in CoT reasoning, Shinn et al. (Shinn et al., 2024) and Madaan et al. (Madaan et al., 2024) employ model to reflect on task feedback signals that can induce better decision-making in subsequent trials. (Zheng et al., 2023a) exploit previously generated answer as hint to progressively guide towards correct answer. Although these methods effectively enhance the performance of LLM, they struggle to produce novel ideas once they have determined a response, as they rely solely on internal representations for generation (Huang et al., 2023). Researchers are currently developing multi-agent collaborative systems to address above issues in pure-textual scenarios (Yin et al., 2023). By designing these systems, large language models (LLMs) can work together to complete tasks or engage in productive debates by offering contrasting perspectives (Liang et al., 2023; Chan et al., 2023; Du et al., 2023). Zhang et al. (Zhang et al., 2023a) further reveal the collaboration mechanism from a social psychology view. This paper represents an initial endeavor to expand upon this method to facilitate multimodal reasoning. By incorporating multiple perspectives from different multimodal language models, we can help address some of the limitations of individual models. Moreover, we address the trivialization of opinions and focus diversion problems of vanilla multi-agent debate via Blueprint Debate on Graph (BDoG).
2.3. Graph-augmented LLMs
Prior research has investigated the integration of structured graphs, such as knowledge graphs (KGs), into large language models (LLMs) by embedding the knowledge into the underlying neural networks (Lin et al., 2019; Wang et al., 2023d). Nevertheless, embedding KGs within LLMs may compromise the inherent explainability and adaptability associated with knowledge reasoning and updating (Hu et al., 2023). To tackle these challenges, recent studies have put forth innovative solutions. Li et al. (Li et al., 2023b) proposed an adaptive query generator, facilitating the creation of queries across various query languages (e.g., SPARQL) to infer rationales. Wang et al. (Wang et al., 2023a) devised a structured multi-round question-answering (QA) format, which extracts external knowledge and generates coherent reasoning traces grounded in precise answers. Sun et al. (Sun et al., 2023) introduced Think-on-Graph (ToG), a method that sequentially reasons over KGs to locate relevant triples, thereby supporting the LLM in predicting the final answer. In the context of multimodal reasoning, CCoT (Mitra et al., 2023) substitutes the rationale generation process with scene graph extraction to enhance the compositional capabilities of large multimodal models. KAM-CoT (Mondal et al., 2024), on the other hand, incorporates external KGs during the two-stage training process, yielding state-of-the-art fine-tuning outcomes in multimodal reasoning. In contrast to existing methods that utilize static graphs, our proposed BDoG preserves the dynamics and precision of KGs through iterative updates of entities, attributes, and relationships, guided by a blueprint debate process.
3. Preliminary
We begin by outlining existing approaches for tackling the multimodal reasoning problem. Figure 2 shows the specific distinct among them. Formally, given a question consisting of tokens, our goal is to identify the correct answer from a set of candidate answers. In the context of multimodal reasoning, the expected answer is intended to be inferred based on a visual context and a textual clue , in addition to the question itself.
Vanilla Prompting. Vanilla prompting approaches aim to predict an answer by augmenting the input with illustrative examples in addition to the question , visual context , and textual clue .
Multimodal CoT. As noted by Lu et al. (Lu et al., 2022), incorporating intermediate reasoning steps (rationales) can aid in predicting the correct answer, especially for complex multimodal reasoning tasks. To address this, we first generate a rationale given the input. The generated rationale is then concatenated with the original language input to update the language representation. This augmented language input is fed together with the original visual input into the same model to infer the final answer.
DDCoT. The Duty-Distinct Chain of Thought framework proposes a novel approach for deconstructing questions into fundamental sub-questions, similar to breaking down reasoning into elementary steps. Contrary to prior work on conversational agents, Zheng et al. (Zheng et al., 2023b) employ the instruction please decompose the question sequentially into necessary sub-inquiries to acquire the sub-question sequence in a single interaction. Within this framework, the final response is obtained by aggregating the answers to each sub-question and the generated CoT rationale .
Self-Correction. Self-correction techniques (Welleck et al., 2022) endeavor to iteratively enhance model predictions by leveraging feedback generated from the model itself. In particular, a feedback function is adopted to iteratively map model outputs to the refined responses.
MAD. Multi-agent Debate (Liang et al., 2023) presents a promising framework that fosters discursive exchange and cross-pollination of ideas between conversational models. Consider a debate comprising rounds amongst a set of large language models acting as interlocutors. In each round, The proponent generates a rationale and response revised in light of the rationales presented by the opponent in prior turns.
4. Blueprint Debate on Graph
In this section, we introduce Blueprint Debate-on-Graph (BDoG). As illustrated in Figure 2, BDoG takes a deductive approach instead of inducing answers from word-level thoughts. It utilizes graphs to structure the opinions and proposals provided by agents. This graph-level structuring of the debating context helps minimize opinion trivialization and focus diversion. Furthermore, BDoG adopts an top-down approach which improves multimodal reasoning by iteratively refining an initial proposal, represented as a blueprint graph. This integrates opinions from diverse perspectives through the competition and cooperation among multiple agents.
The debating progress of BDoG at the round can be formulated as a quadruple
where, given a multimodal source set , the debating is conducted among a set of agents , in which each agent can use operations from the set of to propose/summarize opinions by refining the graph-of-thoughts . At the end of the round, is updated to to initiate the next round.
4.1. Blueprint Initialization
To initiate the debating, we need to convert the multimodal sources into a blueprint graph. This conversion is achieved through the operation function . To implement , we define two additional sub-functions and for extracting entities and relations from the textual sources (i.e., and ) and visual source (i.e., ), respectively. The implementation of is formulated as
where denotes the union of two sets of graphs. The 2 constraints are as follows: 1) Size Constraint: The size of needs to be restricted within a specific range to prevent an excessive number of clues that could distract the inference or an insufficient number to answer the question effectively. 2) Relevance Constraint: We should merge the relationships extracted from and towards those of the question , ensureing all the knowledge encapsulated in is relevant to the question.
Extensive libraries are available for and , as they have been extensively researched (e.g., named entity recognition, relation extraction for , image captioning, visual grounding for ). However, the recent advancements in multimodal large language models (MLLM) have made it convenient to implement these sub-functions using in-context learning based prompts. For example, to extend the query in the context, we can employ CoT to implement as : Given the question {Q}, please provide the necessary steps to answer this question. where the { } denotes the placeholder in the prompt.
For , its implementation varies depending on LLMs used. For GPT-4, the image needs to be encoded in Base64 format. Gemini utilizes PIL for image encoding. InstructBLIP offers its EVA-G encoder to convert the image into an eigenvector. The can then be implemented as : Given the image {} and question {}, generate a scene graph with evidence to answer the question. Please ensure adherence to following constrains: {}, {}. where two exemplar constraints are The graph must not be empty. Please restrict the maximum number of objects in the graph to 20. The objects and relations within the graph should be pertinent to addressing the question.
It worth mentioning that although we provide some exemplar implementations of functions and constraints, the effectiveness of prompts can vary significantly depending on the MLLM used. The success of multimodal reasoning relies more on the development of guiding principles for prompting the models and constraints for regularizing the resulting graph. Therefore, in the rest of this section, our focus lies on discussing these guiding principles and constraints. Our prompt implementations will be provided in Appendix.
4.2. Agents and Roles
In the debate, we can treat each LLM as an agent that participates in the discussion by refining the blueprint graph . Just like in a real debate, each agent has a distinct role assigned. We define three roles as a set of . These roles not only help structure the discussion but also promote critical thinking and ensure a comprehensive and in-depth exploration of the topic.
Proponent agents advocate and defend the current blueprint by refining current into an affirmative evidence graph . A debating function is assigned for this purpose as
An exemplar implementation is : As {}, you are assigned as an affirmative debater and have been provided with an evidence graph {} for answering the question {} related to the image {}. Try to enhance the graph by incorporating your insights towards an optimal solution. Please ensure adherence to following constrains: {}, {}, {}. Note that we have incorporated the conclusion from (Du et al., 2023; Yin et al., 2023) that the agent’s understanding of the role can be improved by using the {} for targeted personality injection. Furthermore, the personality can be tailored to be specific, such as “Ben, a high school student with an impressive academic record and respected by peers for your knowledge and logical thinking.” The Proponent debate adheres to the Size and Relevance constraints defined in Eq. (3), and it also includes the Compactness Constraint: The refined graph should be as concise as possible, ensuring that the blueprint remains focused.
Opponent agents challenge and present arguments against the blueprint by updating it into a negative evidence graph as
An exemplar implementation is : As {}, you are assigned as a negative debater and have been provided with an affirmative evidence graph {} for answering the question {} regarding the image {}. Try to detect potential flaws and drawbacks of the graph and update it with your insights. Please ensure adherence to following constrains: {}, {}, {}. The utilization of the functions and fosters an adversarial dynamic between the Proponent and Opponent, ensuring a diverse and comprehensive discussion.
To facilitate the debating, Moderator agents synthesizes the arguments and opinions presented by both the proponent and opponent by merging the and into a conclusion graph as
An exemplar implementation is : As {}, you are assigned as a moderator in a debate and have been provided with an affirmative evidence graph {} and a negative evidence graph {} to address the question {} regarding the image {}. Try to consolidate the two graphs into a single graph towards the optimal solution, and provide a conclusive answer to the question.
4.3. Debate Progress and Graph Condensation
Initialization and Role Assignment: Once the blueprint has been initialized, the debate commences with the assignment of roles to agents in . Denote the assignment of a role to an agent as , to ensure a balanced debate, an equal number of agents are assigned as Proponents and Opponents, with only one agent assigned as the Moderator. The Role Assignment Regulation is written
Debating: After roles are assigned, the debate can be conducted iteratively between the Proponents and Opponents as illustrated in Figure 2. The initial blueprint is then updated in subsequent debate rounds. In each round, the Moderator summarizes the affirmative and negative graphs in a conclusion graph on the basis of which a tentative answer is also provided. If the debate is not concluded, the Moderator initiates the next round by assign as the blueprint . Otherwise, the Moderator’s answer is considered final and adopted.
Stopping Criteria: The condition to conclude the debate can be determined by assessing the modifications made to the evidence graph compared to the previous round as
where is a distance metric defined on the graphs. The rationale is that with each successful round of debate, the evidence becomes more concise, leading to the condensation of the evidence graph. Therefore, we can quantify the modification by tallying the number of entities (relations) that have been updated and pruned as
5. Experiments
5.1. Backbone Models
To evaluate its performance and generalizability, we have implemented Blueprint Debate-on-Graph (BDoG) using different prevalent multimodal large language models as backbones, including 1) GeminiProVision (Team et al., 2023), an extensively parameterized model developed by Google, 2) InstructBLIP (Dai et al., 2023), which possesses more constrained dimensions and computational resources relative to alternative architectures, and 3) GPT-4 which is the fourth iteration of the GPT model developed by OpenAI. More implementation details can be found from the Appendix.
5.2. Datasets and Metrics
In line with the general setup described in (Zheng et al., 2023b; Lu et al., 2024), we perform our experiments using two extensively adopted multimodal question answering (QA) datasets. These datasets are widely recognized as standard benchmarks, specifically designed to evaluate the performance and effectiveness of models in addressing multimodal reasoning tasks. The two benchmarks are: 1) ScienceQA-IMG (SQA-IMG) (Lu et al., 2022) represents the first multimodal scientific question-answering corpus comprising 21,000 inquiries paired with multiple choices and accompanying images. As a training-free approach, we solely utilize the TEST and DEV partitions of ScienceQA-IMG following prior work (Lu et al., 2022) for comparative assessment. Additionally, the questions within ScienceQA can be categorized into three subdomains: natural sciences (NAT), social sciences (SOC), and linguistic sciences (LAN). 2) MMbench (Liu et al., 2023) offers a more systematic and robust means for reasoning evaluation compared to existing benchmarks such as VQAv2 or COCO Captions. We employ the official data split (MMBench-Dev) and code released by the originating authors. The statistics of the two benchmarks are delineated in Appendix. We report the accuracy metric as determined through a heuristic matching procedure, following the same setting of the official benchmark (Lu et al., 2022).
| Model | Size | SQA-IMG | MMBench |
| MiniGPT-4 (Zhu et al., 2024) | 7B | 37.7 | 24.3 |
| Qwen-VL (Bai et al., 2023) | 7B | 58.6 (67.1) | 38.2 |
| Qwen-VL-Chat (Bai et al., 2023) | 7B | 68.6 (68.2) | 60.6 |
| mPLUG-Owl2 (Xu et al., 2023) | 8B | 63.9 | 66.5 |
| CogVLM-Chat (Wang et al., 2023c) | 17B | 69.6 | 63.7 |
| LLaVA-v1.5 (Liu et al., 2024) | 13B | 71.9 (71.6) | 68.2 |
| InstructBLIP (Dai et al., 2023) | 13B | 59.2 (63.1) | 36.0 |
| InstructBLIP+BDoG | 13B | 63.5 | 55.8 |
| GPT-3.5+CoT (Wei et al., 2022) | 175B | 67.4 | - |
| GPT-3.5+DDCoT (Zheng et al., 2023b) | 175B | 72.5 | - |
| GPT-4+CoT (Wei et al., 2022) | - | 71.5 | 75.1 |
| GPT-4+BDoG | - | 77.2 | 79.2 |
| GeminiProVision (Team et al., 2023) | - | 76.5 | 75.2 |
| GeminiProVision+BDoG | - | 81.1 | 81.3 |
5.3. Performance Comparison to SOTA Methods
We evaluate the proposed method by by comparing it against two sets of state-of-the-art approaches as follows:
Open-Source Multimodal LLMs with Relatively Moderate Parameters including MiniGPT-4 (Zhu et al., 2024), Qwen-VL and Qwen-VL- Chat (Bai et al., 2023), CogVLM-Chat (Wang et al., 2023c), mPLUG-Owl2 (Xu et al., 2023), LLaVA-v1.5 (Liu et al., 2024), and InstructBLIP (Dai et al., 2023). These models have parameter scales ranging from 7B to 17B
Closed-Source Multimodal LLMs with Large-Scale Parameters: GPT-3.5 (Wei et al., 2022), GPT-4V (Achiam et al., 2023) and GeminiProVision (Team et al., 2023). Following the general standard, GPT-3.5 and GPT-4 have been incorporated with the CoT (Wei et al., 2022) or DDCoT (Zheng et al., 2023b) (built based on image captioning results). These models are known for their parameter scales above 175B and are considered to have the best performance in most of the literature.
The results are shown in Table 1. The integration of BDoG has resulted in a significant improvement across different backbones, as evidenced by the performance gains of on SQA-IMG and on MMBench. Notably, when combined with GeminiProVision, BDoG achieves SOTA performance on the ScienceQA-IMG test set and MMBench development set, achieving accuracies of 81.1% and 81.3%, respectively. Other observations that indicate BDoG’s advantage over SOTA methods include:
BDoG helps reduce the performance gap between large and small models. It is commonly believed that models with larger parameter scales tend to perform better than smaller ones. This observation generally holds true, as shown in Table 1 for models without BDoG. However, the introduction of BDoG has led to a reduction in the performance gap between these two types of models. This can be seen in the improvement achieved by InstructBLIP, which has experienced a boost of 4.3% and achieves an accuracy of 63.5% on SQA-IMG, comparable to that of GPT-3.5.
BDoG reinforces the multimodal reasoning. Form Table 1, we can also observe the advantage of direct multimodal reasoning (e.g., open-source VL models, and GeminiProVision) over indirect multimodal reasoning (e.g., GPT3.5+CoT and GPT3.5+DDCoT due to their nature of obtaining visual information through image captioning). Even the open-source VL models of the former group achieves comparable performance to those of the latter one, with much smaller parameter scales. With BDoG, which reinforces multimodal reasoning by graph regulation, the performance of direct multimodal reasoning of InstructBLIP and GeminiProVision have been improved by 6.1% and 19.8% on the MMBench dataset.
| Model | Method | ScienceQA-IMG-Dev | ScienceQA-IMG-Test | MMBench-Dev | ||||||||||||
| NAT | SOC | LAN | Avg | NAT | SOC | LAN | Avg | LR | AR | RR | FP-S | FP-C | CP | Avg | ||
| MniGPT-4 (Zhu et al., 2024) | Base | 42.9 | 30.6 | 43.7 | 38.4 | 42.0 | 30.1 | 50.0 | 37.7 | 7.5 | 31.3 | 4.3 | 30.3 | 9.0 | 35.6 | 24.3 |
| Qwen-VL (Bai et al., 2023) | 52.1 | 59.8 | 58.3 | 55.0 | 55.7 | 62.0 | 77.3 | 58.7 | 16.1 | 44.7 | 34.8 | 35.2 | 39.2 | 46.6 | 38.2 | |
| Qwen-VL-Chat (Bai et al., 2023) | 60.9 | 67.4 | 62.5 | 63.3 | 67.7 | 69.6 | 75.0 | 68.6 | 32.2 | 59.8 | 43.5 | 66.2 | 48.3 | 79.4 | 60.6 | |
| mPLUG-Owl2 (Xu et al., 2023) | 60.6 | 68.0 | 45.8 | 62.8 | 62.5 | 66.2 | 61.4 | 63.9 | 32.2 | 72.4 | 60.9 | 68.6 | 60.1 | 79.4 | 66.5 | |
| CogVLM-Chat (Wang et al., 2023c) | 63.1 | 69.2 | 77.1 | 65.6 | 68.0 | 72.2 | 70.4 | 69.7 | 29.7 | 65.8 | 60 | 66.9 | 58 | 76.7 | 63.7 | |
| LLaVA-v1.5 (Liu et al., 2024) | 66.1 | 74.9 | 72.9 | 69.4 | 70.1 | 74.2 | 81.8 | 71.9 | 44.1 | 67.3 | 60.0 | 72.0 | 59.4 | 82.1 | 68.2 | |
| InstructBLIP (Dai et al., 2023) | Base | 53.7 | 57.3 | 47.9 | 54.8 | 58.1 | 61.0 | 61.4 | 59.2 | 14.2 | 46.3 | 22.6 | 37.0 | 21.4 | 49.0 | 36.0 |
| + BDoG | 59.7 | 55.6 | 54.2 | 58.1 | 63.1 | 58.2 | 72.7 | 61.4 | 43.7 | 57.8 | 31.5 | 42.7 | 19.5 | 39.1 | 42.6 | |
| + BDoG | 58.1 | 61.3 | 52.1 | 59.0 | 60.6 | 62.6 | 68.2 | 61.5 | 58.8 | 65.5 | 41.2 | 51.2 | 18.6 | 46.1 | 51.1 | |
| + BDoG | 61.1 | 64.0 | 52.1 | 61.9 | 61.1 | 66.5 | 75.0 | 63.5 | 63.3 | 71.9 | 37.8 | 56.3 | 20.3 | 59.1 | 55.8 | |
| GeminiProVision (Team et al., 2023) | Base | 68.9 | 81.6 | 75.0 | 73.7 | 72.9 | 81.5 | 88.6 | 76.5 | 55.9 | 80.4 | 73.9 | 79.5 | 61.5 | 82.1 | 75.2 |
| + BDoG | 73.3 | 81.1 | 77.1 | 76.2 | 75.3 | 82.8 | 93.2 | 78.5 | 71.1 | 85.1 | 83.1 | 78.9 | 71.9 | 81.3 | 79.3 | |
| + BDoG | 69.8 | 84.8 | 87.5 | 75.6 | 74.7 | 86.8 | 88.6 | 79.6 | 75.0 | 84.5 | 80.7 | 81.4 | 73.0 | 83.6 | 80.7 | |
| + BDoG | 73.6 | 86.2 | 85.4 | 78.4 | 76.6 | 87.4 | 93.2 | 81.1 | 74.0 | 84.8 | 83.4 | 81.3 | 73.7 | 84.4 | 81.3 | |
5.4. Ablation Study
In order to gain a comprehensive understanding of BDoG, we conduct an ablation study where we decompose BDoG into two variants:
BDoG: we remove the graph regulation and constraints, resulting in a debate-only approach (i.e., vanilla multi-agent debate) for investigating the specific contribution of the debating component of BDoG.
BDoG: we remove the debating rounds, resulting in a graph-based reasoning method for investigating the specific contribution of the graph regulation component of BDoG.
Moreover, we analyze the performance of the two variants on the benchmarks by breaking it down into subcategories. This analysis allows us to investigate the preferences of these two variants for different types of questions. The results are presented in Table 2, where it can be observed that both variants demonstrate comparable performance across various benchmarks. This suggests that the debate and graph components of BDoG contribute to its effectiveness in a similar manner. Through the combination of these two components in BDoG, the performance has experienced further improvement compared to the individual variants. However, when considering specific categories, distinctions in the contributions of the debate and graph components become apparent.
Impact of the debate component: BDoG demonstrates consistent improvements across both benchmarks with a debate-only setting, which encourages LLM agents to collaboratively refine and correct prior responses. For science questions, BDoG facilitates the model’s focus on specific errors, such as direction, size, and position, leading to improved performance in the natural science domain (boosting accuracy from 53.7 to 59.7 for InstructBLIP and 68.9 to 73.3 for GeminiProVision). Additionally, BDoG significantly enhances logical reasoning tasks (improving accuracy from 14.2 to 43.7). However, the debate-only nature has limitations, including trivialization and focus diversion issues. Without the graph regulation, overall performance decreases from 55.8 to 42.6 for InstructBLIP, particularly when addressing questions that require attention to specific attributes and relations.
Impact of the graph regulation: With a graph-regularized knowledge base for the discussion, BDoG also demonstrates consistent improvement of overs the base models on both benchmarks. Compared to the text-based and debate-only method BDoG, it performs evidently better on the MMBench by . This is not a surprise, because MMBench emphasizes more the attribute and relations which challenges a model’s ability on structural reasoning. Although incorporating fact-related graph information proves beneficial in BDoG, the absence of the iteratively refined debate procedure results in decreased performance due to the coarse and distorted extraction of blueprint information.
Impact of combining the debate and graph components: By combining the two components, BDoG achieves gains across nearly all categories. In the ScienceQA-IMG dataset, BDoG exhibits consistent and steady improvements, averaging around 5% compared to the baseline models. This suggests that BDoG is robust and generalizes well for science-related questions. Remarkably, BDoG significantly outperforms the baseline model (InstructBLIP) on the MMBench-Dev set, particularly in the areas of Logical Reasoning (LR) with a margin of 49.1%, Attribute Reasoning (AR) with a margin of 26.6%, and Relation Reasoning (RR) with a margin of 15.2%. BDoG enhances logical reasoning (LR) through a mechanism that refines the reasoning process iteratively, emphasizing the importance of multi-step reasoning rationales. The blueprint graph structure of BDoG, which explicitly models objects, attributes, and relations, contributes to improved reasoning abilities in Attribute Reasoning (AR) and Relation Reasoning (RR). The GeminiProVision model also exhibits comparable performance improvements, with BoG contributing to enhanced fine-grained perception across instances (FP-C), resulting in a gain of 12.2%. This improvement can be attributed to the connections established between various objects within the debate-on-graph framework.
A case study for the iterative improvement on the blueprint: BDoG leverages the advantages of both structured evidence through graph regulation and iterative refinement through debating. This is evident in the consistent improvement observed on the blueprint graph, showcasing the combined benefits of these two components. Figure 3 provides running examples demonstrating the superior reasoning performance of our proposed BDoG framework compared to the BDoG method.
The left case draws from the ScienceQA dataset, testing geographic knowledge and map interpretation. While BDoG correctly answered Dominica is highlighted, it also generated irrelevant information about Dominica’s economic development. This misguided the agents into off-topic discussion, concluding incorrectly with Barbados. In contrast, BDoG concentrated on the question and options, iteratively refining the blueprint entities and relations to arrive at the right answer of Dominica.
The example on the right comes from the MMBench dataset requiring cross-instance perception. As the image contained both candies and jars, it posed a challenge. With BDoG relying on text alone, agreement was rarely reached as responses changed over debate rounds. However, BDoG first generated a blueprint defining image objects and attributes. This established the discussion scope. BDoG then pruned irrelevant candy information, focusing discussion on the specific object - jars. It output the final answer by comparing and connecting the two jar sub-graphs.
In summary, Figure 3 demonstrates that BDoG beats BDoG on both datasets through its blueprint-driven approach. This concentrates graph-based reasoning on salient topics and prunes irrelevant details to arrive at well-supported conclusions.
| Round | ScienceQA-IMG-Test | MMBench-Dev | ||
| BDoG-S | BDoG-L | BDoG-S | BDoG-L | |
| 1 | 60.5 | 80.6 | 51.6 | 81.0 |
| 2 | 63.5 | 80.9 | 54.6 | 81.1 |
| 3 | 63.1 | 81.1 | 55.8 | 81.3 |
| 4 | 63.3 | 81.4 | 55.8 | 80.9 |
0pt
olFFL
5.5. Monitoring The Debating Progress
We evaluated the model’s performance against the termination criteria across multiple debate rounds based on the data in Table 3. Our analysis showed that for models with smaller parameters like InstructBLIP, moving from a single round to two rounds led to significant gains in performance. This improvement was particularly notable when increasing the number of rounds from one to two. However, for larger models that may reach agreement more easily, the performance enhancement was relatively modest when amplifying the number of debate rounds. In general, we found the model’s performance tended to converge within the second or third round. This can be attributed to the underlying reasoning typically being able to answer questions within 2-3 steps.
Additionally, Figure 4 illustrates the number of updated attributes, newly added or removed entities or relations between and within rounds. A strength of our proposed BDoG framework is its ability to quantify the debate process by inspecting graph changes. This demonstrates the effectiveness of dynamically adjusting the initial graph based on the discussion. The results in Figure 4 are also consistent with our hypothesis that disagreements and errors can be decreased as the debate progresses.
5.6. Efficiency Analysis
We further compared the effectiveness versus efficiency of our BDoG framework against BDoG, with results shown in Figure 5. Maintaining concise content focused on key aspects, the graph structure of BDoG demonstrated superior efficiency, requiring approximately 50% less inference time than BDoG. By first generating a blueprint, BDoG defines the scope of the current state, thereby improving model efficiency by filtering irrelevant information. Concurrently, Figure 5 shows BDoG outperformed BDoG in effectiveness, achieving over 5 percentage points higher accuracy than BDoG across three test sets.
This enhanced effectiveness can be attributed to BDoG’s framework concentrating on important knowledge rather than generational textual content without guidance, as in BDoG. Defining a blueprint to establish the discussion boundaries before generating a graph structure enhances both efficiency and effectiveness by restricting the model’s reasoning to salient topics. In summary, Figure 5 illustrates BDoG demonstrated better time performance and predictive power compared to BDoG, validating the utility of our approach in balancing computational cost and predictive accuracy through focused, blueprint-driven debate.
6. Conclusion
This paper has presented a pioneering pilot study that introduces multi-agent debate into the realm of multimodal reasoning. We tackled two prominent challenges faced in this context: the issue of opinions being trivialized and focus diversion. By recognizing the limitations of existing debating schemes, we propose Blueprint Debate on Graphs (BDoG), which confines debates to a blueprint graph and stores evidence in graph branches, to address the challenges of word-level opinion trivialization and distraction caused by irrelevant concepts. Extensive experiments conducted in Science QA and MMBench validate the efficacy of BDoG, surpassing previous methods and establishing new state-of-the-art results.

### v2

@@
-A Picture Is Worth a Graph: Blueprint Debate on Graph for Multimodal Reasoning
+A Picture Is Worth a Graph: A Blueprint Debate Paradigm for Multimodal Reasoning
 Abstract.
@@
 Moreover, by storing evidence in branches within the graph, BDoG mitigates distractions caused by frequent but irrelevant concepts.
-Extensive experiments validate BDoG, achieving state-of-the-art results in Science QA and MMBench with significant improvements over previous methods.
+Extensive experiments validate that BDoG is able to achieve state-of-the-art results in ScienceQA and MMBench with significant improvements over previous methods.
+The source code can be accessed at https://github.com/thecharm/BDoG.
 1. Introduction
 Multimodal reasoning depends on two key aspects: the creation of a unified representation of semantics from different modalities and the integration of these diverse semantics while ensuring logical consistency.
-While the advancement in large language models (LLMs) has made it possible to represent the semantics in natural languages (Achiam et al., 2023), the integration of diverse semantics remains a challenging issue, even in exclusive NLP tasks.
+While the advancement in large language models (LLMs) has made it possible to represent the semantics in natural languages (Achiam et al., 2023; Team et al., 2023), the integration of diverse semantics remains a challenging issue, even in exclusive NLP tasks.
 One approach to tackle this challenge is multi-agent debate (MAD), where multiple LLMs act as agents, each contributing their own perspectives on the target topic and reaching a consensus through debates (Liang et al., 2023; Chan et al., 2023).
 This scheme could be adopted by incorporating a specific LLM for each modality as an agent.
-Despite being relatively unexplored in the multimodal domain, MAD encounters numerous challenges in a broader context.
-MAD may suffer from the trivialization of opinions, which is a result of the summarization step performed at the conclusion of each debating round.
+While being relatively unexplored in the multimodal domain, MAD encounters numerous challenges in a broader context.
+It may suffer from the trivialization of opinions, resulting from the summarization step performed at the conclusion of each debating round.
 The objective of this step is to seek agreement among the participating agents regarding their opinions.
 Consequently, this process can lead to the debate’s focus being directed towards a general concept, serving as an adaptation to accommodate the diverse range of semantics.
-One example can be observed in the reasoning of the Multimodal Language Model (MLLM) depicted in Figure 1, where the image modality presents a diverse range of semantics, including bear sedge, earthworm, collared lemming, and others.
+One example can be observed in the reasoning of the Multimodal Large Language Model (MLLM) depicted in Figure 1, where the image modality presents a diverse range of semantics, including bear sedge, earthworm, collared lemming, and others.
 As a consequence, this can result in the context and summary being trivialized, shifting the emphasis from lichen to a more generalized concept of the tundra ecosystem, wherein both bilberry and mushroom exhibit a high degree of correlation.
@@
 The semantic divergence increases further when the impacts of those concepts are amplified through CoT, particularly when the newly introduced concepts exhibit biases towards certain concepts, resulting in focus diversion.
-To address this issue, we propose an deductive reasoning scheme called Blueprint Debate on Graph (BDoG, pronounced bee-dog).
-BDoG is inspired by the blueprint debate that has been widely employed in real-world debates, which distinguishes itself from other debates by its concentration on evaluating and refining a proposal (e.g., blueprint) to address specific challenges or issues.
-BDoG begins by aggregating concepts from modalities and incorporating with their relationships into a initial graph.
+To address these issues, we propose an deductive reasoning scheme called Blueprint Debate on Graph (BDoG, pronounced bee-dog).
+BDoG is inspired by the blueprint debate that has been employed in real-world debates, which distinguishes itself from other debates by its concentration on evaluating and refining a proposal (e.g., blueprint) to address specific challenges or issues.
+BDoG begins by aggregating concepts from modalities and incorporating with their relationships into an initial graph.
 This graph serves as a blueprint that confines the scope of the discussion rather than having it open to irrelevant semantics as in existing schemes.
@@
 The newly introduced concepts are incorporated into relevant branches instead of remaining as a word-level thoughts within the context.
-This reduces the likelihood of focus diversion since, in DBoG, the competition of semantics occurs at the branch level rather than the word level.
-This can be seen from Figure 1, where the most relevant branches related to the soil and caribou standout from the competition, eliminating the irrelevant semantcis effectively.
-In addition to the advantages of scope-confined guidance and branch-level competition, BDoG also increases explainability, allowing for the tracking of discussion progress.
-This is evident in Figure 1, in which the progress of debating is more understandable than word-level reasoning (i.e., MLLM-CoT or MAD).
+This reduces the likelihood of focus diversion since, in BDoG, the competition of semantics occurs at the branch level rather than the word level.
+This can be seen from Figure 1, where the most relevant branches related to the soil and caribou standout from the competition, eliminating the irrelevant semantics effectively.
+In addition to the advantages of scope-confined guidance and branch-level competition, BDoG also increases explainability, allowing for the tracking of discussion progress (Figure 1).
 2. Related Works
 2.1. Multimodal Reasoning
-Multimodal reasoning is a crucial component in the development of advanced artificial intelligence (AI) systems that aim to replicate human-like intelligence (Lu et al., 2022).
-This type of reasoning enables AI systems to process and analyze information from various sources and forms, such as text, images, audio, and video, in a integrated and coordinated manner (Nam et al., 2017; Cadene et al., 2019).
-The latest advancements in multimodal large language models, such as BLIP2 (Li et al., 2023a), KOSMOS (Huang et al., 2024) and LLaVA (Liu et al., 2024) have demonstrated significant progress in complex reasoning, as these models (Zhang et al., 2023b) now have the capability to generate step-by-step rationales prior to producing the final answer, following a chain-of-thought (CoT) manner.
+Multimodal reasoning is a crucial component in the development of advanced artificial intelligence (AI) systems that aim to replicate human-like intelligence (Lu et al., 2022; lv2023duet).
+The latest advancements in multimodal large language models, such as BLIP2 (Li et al., 2023b) and LLaVA (Liu et al., 2024) have demonstrated significant progress in complex reasoning, as these models (Zhang et al., 2023b) now have the capability to generate step-by-step rationales prior to producing the final answer, following a chain-of-thought (CoT) manner.
 Zheng et al.
@@
 Chameleon (Lu et al., 2024) accomplishes complex multimodal reasoning tasks by integrating various external tools (e.g., large language models, off-the-shelf vision models, and web search engines).
-Nevertheless, existing methods exhibit limitations as they heavily rely on either few-shot learning or supervision to “guide” the reasoning process.
-In order to overcome this dependency, we propose the incorporation of debating features into our method.
-This enables the agents to engage in an adversarial discussion, allowing them to “figure out” the correct direction autonomously.
-As a result, our approach makes zero-shot learning feasible by reducing the reliance on external guidance or supervision.
 2.2. Multi-agent Debate
@@
 2.3. Graph-augmented LLMs
-Prior research has investigated the integration of structured graphs, such as knowledge graphs (KGs), into large language models (LLMs) by embedding the knowledge into the underlying neural networks (Lin et al., 2019; Wang et al., 2023d).
+Prior research has investigated the integration of structured graphs, such as knowledge graphs (KGs), into large language models (LLMs) by embedding the knowledge into the underlying neural networks (wei2008fusing; Wang et al., 2023d).
 Nevertheless, embedding KGs within LLMs may compromise the inherent explainability and adaptability associated with knowledge reasoning and updating (Hu et al., 2023).
@@
 Li et al.
-(Li et al., 2023b) proposed an adaptive query generator, facilitating the creation of queries across various query languages (e.g., SPARQL) to infer rationales.
+(Li et al., 2023c) propose an adaptive query generator, facilitating the creation of queries across various query languages (e.g., SPARQL) to infer rationales.
 Wang et al.
-(Wang et al., 2023a) devised a structured multi-round question-answering (QA) format, which extracts external knowledge and generates coherent reasoning traces grounded in precise answers.
+(Wang et al., 2023a) devise a structured multi-round question-answering (QA) format, which extracts external knowledge and generates coherent reasoning traces grounded in precise answers.
 Sun et al.
-(Sun et al., 2023) introduced Think-on-Graph (ToG), a method that sequentially reasons over KGs to locate relevant triples, thereby supporting the LLM in predicting the final answer.
+(Sun et al., 2023) introduce Think-on-Graph (ToG), a method that sequentially reasons over KGs to locate relevant triples, thereby supporting the LLM in predicting the final answer.
 In the context of multimodal reasoning, CCoT (Mitra et al., 2023) substitutes the rationale generation process with scene graph extraction to enhance the compositional capabilities of large multimodal models.
@@
 We begin by outlining existing approaches for tackling the multimodal reasoning problem.
-Figure 2 shows the specific distinct among them.
+Figure 2 shows the specific distinction among them.
 Formally, given a question consisting of tokens, our goal is to identify the correct answer from a set of candidate answers.
@@
 Contrary to prior work on conversational agents, Zheng et al.
-(Zheng et al., 2023b) employ the instruction please decompose the question sequentially into necessary sub-inquiries to acquire the sub-question sequence in a single interaction.
+(Zheng et al., 2023b) employ the instruction to acquire the sub-question sequence in a single interaction.
 Within this framework, the final response is obtained by aggregating the answers to each sub-question and the generated CoT rationale .
@@
 MAD.
-Multi-agent Debate (Liang et al., 2023) presents a promising framework that fosters discursive exchange and cross-pollination of ideas between conversational models.
-Consider a debate comprising rounds amongst a set of large language models acting as interlocutors.
-In each round, The proponent generates a rationale and response revised in light of the rationales presented by the opponent in prior turns.
+MAD (Liang et al., 2023) presents a promising framework that fosters discursive exchange and cross-pollination of ideas between conversational models.
+Consider a debate comprising rounds amongst a set of large language models acting as interlocutors, the proponent generates a rationale and response revised in the light of rationales presented by the opponent in prior turns.
 4. Blueprint Debate on Graph
@@
 It utilizes graphs to structure the opinions and proposals provided by agents.
-This graph-level structuring of the debating context helps minimize opinion trivialization and focus diversion.
-Furthermore, BDoG adopts an top-down approach which improves multimodal reasoning by iteratively refining an initial proposal, represented as a blueprint graph.
+This graph-level structuring of the debating context helps to minimize opinion trivialization and focus diversion.
+Furthermore, BDoG adopts a top-down approach which improves multimodal reasoning by iteratively refining an initial proposal, represented as a blueprint graph.
 This integrates opinions from diverse perspectives through the competition and cooperation among multiple agents.
-The debating progress of BDoG at the round can be formulated as a quadruple
-where, given a multimodal source set , the debating is conducted among a set of agents , in which each agent can use operations from the set of to propose/summarize opinions by refining the graph-of-thoughts .
+The BDoG at the round can be formulated as a quadruple
+where, given a multimodal source set , the debating is conducted among a set of agents , in which each agent uses operations from the set to propose opinions by refining the graph-of-thoughts .
 At the end of the round, is updated to to initiate the next round.
@@
 The 2 constraints are as follows: 1) Size Constraint: The size of needs to be restricted within a specific range to prevent an excessive number of clues that could distract the inference or an insufficient number to answer the question effectively.
-2) Relevance Constraint: We should merge the relationships extracted from and towards those of the question , ensureing all the knowledge encapsulated in is relevant to the question.
-Extensive libraries are available for and , as they have been extensively researched (e.g., named entity recognition, relation extraction for , image captioning, visual grounding for ).
+2) Relevance Constraint: We should merge the relationships extracted from and towards those of the question , ensuring all the knowledge encapsulated in is relevant to the question.
+Extensive libraries are available for and , as they have been extensively researched (e.g., named entity recognition (Wu et al., 2020), relation extraction (Zheng et al., 2021) for , image captioning (Zhao and Wu, 2023), visual grounding (Li et al., 2023a) for ).
 However, the recent advancements in multimodal large language models (MLLM) have made it convenient to implement these sub-functions using in-context learning based prompts.
-For example, to extend the query in the context, we can employ CoT to implement as : Given the question {Q}, please provide the necessary steps to answer this question.
-where the { } denotes the placeholder in the prompt.
+For example, to extend the query in the context, we can employ CoT to implement as {mdframed}[hidealllines=true,backgroundcolor=gray!20]
+: Given the question {Q}, please provide the necessary steps to answer this question. where the { } denotes the placeholder in the prompt.
 For , its implementation varies depending on LLMs used.
@@
 InstructBLIP offers its EVA-G encoder to convert the image into an eigenvector.
-The can then be implemented as : Given the image {} and question {}, generate a scene graph with evidence to answer the question.
+The can then be implemented as {mdframed}[hidealllines=true,backgroundcolor=gray!20]
+: Given the image {} and question {}, generate a scene graph with evidence to answer the question.
 Please ensure adherence to following constrains: {}, {}.
-where two exemplar constraints are The graph must not be empty.
-Please restrict the maximum number of objects in the graph to 20.
+where two exemplar constraints are {mdframed}[hidealllines=true,backgroundcolor=gray!20]
+The graph must not be empty. Please restrict the maximum number of objects in the graph to 20. {mdframed}[hidealllines=true,backgroundcolor=gray!20]
 The objects and relations within the graph should be pertinent to addressing the question.
@@
 Proponent agents advocate and defend the current blueprint by refining current into an affirmative evidence graph . A debating function is assigned for this purpose as
-An exemplar implementation is : As {}, you are assigned as an affirmative debater and have been provided with an evidence graph {} for answering the question {} related to the image {}.
+An exemplar implementation is {mdframed}[hidealllines=true,backgroundcolor=gray!20]
+: As {}, you are assigned as an affirmative debater and have been provided with an evidence graph {} for answering the question {} related to the image {}.
 Try to enhance the graph by incorporating your insights towards an optimal solution.
@@
 Opponent agents challenge and present arguments against the blueprint by updating it into a negative evidence graph as
-An exemplar implementation is : As {}, you are assigned as a negative debater and have been provided with an affirmative evidence graph {} for answering the question {} regarding the image {}.
+An exemplar implementation is {mdframed}[hidealllines=true,backgroundcolor=gray!20]
+: As {}, you are assigned as a negative debater and have been provided with an affirmative evidence graph {} for answering the question {} regarding the image {}.
 Try to detect potential flaws and drawbacks of the graph and update it with your insights.
@@
 The utilization of the functions and fosters an adversarial dynamic between the Proponent and Opponent, ensuring a diverse and comprehensive discussion.
-To facilitate the debating, Moderator agents synthesizes the arguments and opinions presented by both the proponent and opponent by merging the and into a conclusion graph as
-An exemplar implementation is : As {}, you are assigned as a moderator in a debate and have been provided with an affirmative evidence graph {} and a negative evidence graph {} to address the question {} regarding the image {}.
+To facilitate the debating, Moderator agents synthesize the arguments and opinions presented by both the proponent and opponent by merging the and into a conclusion as
+An exemplar implementation is {mdframed}[hidealllines=true,backgroundcolor=gray!20]
+: As {}, you are assigned as a moderator in a debate and have been provided with an affirmative evidence graph {} and a negative evidence graph {} to address the question {} regarding the image {}.
 Try to consolidate the two graphs into a single graph towards the optimal solution, and provide a conclusive answer to the question.
@@
 Denote the assignment of a role to an agent as , to ensure a balanced debate, an equal number of agents are assigned as Proponents and Opponents, with only one agent assigned as the Moderator.
-The Role Assignment Regulation is written
+The Role Assignment Regulation is
 Debating: After roles are assigned, the debate can be conducted iteratively between the Proponents and Opponents as illustrated in Figure 2.
@@
 5.1. Backbone Models
-To evaluate its performance and generalizability, we have implemented Blueprint Debate-on-Graph (BDoG) using different prevalent multimodal large language models as backbones, including 1) GeminiProVision (Team et al., 2023), an extensively parameterized model developed by Google, 2) InstructBLIP (Dai et al., 2023), which possesses more constrained dimensions and computational resources relative to alternative architectures, and 3) GPT-4 which is the fourth iteration of the GPT model developed by OpenAI.
+To evaluate its performance and generalizability, we have implemented Blueprint Debate-on-Graph (BDoG) using different prevalent multimodal large language models as backbones, including 1) GeminiProVision (Team et al., 2023), an extensively parameterized model developed by Google, 2) InstructBLIP (Dai et al., 2023) and LLaVA-v1.5 (Liu et al., 2024), which possesses more constrained dimensions and computational resources relative to alternative architectures, and 3) GPT-4 (Achiam et al., 2023) which is the fourth iteration of the GPT model developed by OpenAI.
 More implementation details can be found from the Appendix.
@@
 As a training-free approach, we solely utilize the TEST and DEV partitions of ScienceQA-IMG following prior work (Lu et al., 2022) for comparative assessment.
-Additionally, the questions within ScienceQA can be categorized into three subdomains: natural sciences (NAT), social sciences (SOC), and linguistic sciences (LAN).
-2) MMbench (Liu et al., 2023) offers a more systematic and robust means for reasoning evaluation compared to existing benchmarks such as VQAv2 or COCO Captions.
+2) MMbench (Liu et al., 2023) offers a more systematic and robust means for zero-shot reasoning evaluation compared to existing benchmarks such as VQAv2 (Goyal et al., 2017) or COCO Captions (Chen et al., 2015).
 We employ the official data split (MMBench-Dev) and code released by the originating authors.
-The statistics of the two benchmarks are delineated in Appendix.
-We report the accuracy metric as determined through a heuristic matching procedure, following the same setting of the official benchmark (Lu et al., 2022).
+We report the accuracy metric through a heuristic matching procedure, following the same setting of the official benchmark (Lu et al., 2022).
+The statistics of the two benchmarks and detailed settings are delineated in Appendix.
 | Model | Size | SQA-IMG | MMBench |
@@
 | CogVLM-Chat (Wang et al., 2023c) | 17B | 69.6 | 63.7 |
-| LLaVA-v1.5 (Liu et al., 2024) | 13B | 71.9 (71.6) | 68.2 |
-| InstructBLIP (Dai et al., 2023) | 13B | 59.2 (63.1) | 36.0 |
+| InstructBLIP (Dai et al., 2023) | 13B | 59.2 (63.1) | 44.0 |
 | InstructBLIP+BDoG | 13B | 63.5 | 55.8 |
+| LLaVA-v1.5 (Liu et al., 2024) | 13B | 71.6 | 68.2 |
+| LLaVA-v1.5+BDoG | 13B | 72.0 | 71.1 |
 | GPT-3.5+CoT (Wei et al., 2022) | 175B | 67.4 | - |
 | GPT-3.5+DDCoT (Zheng et al., 2023b) | 175B | 72.5 | - |
-| GPT-4+CoT (Wei et al., 2022) | - | 71.5 | 75.1 |
-| GPT-4+BDoG | - | 77.2 | 79.2 |
-| GeminiProVision (Team et al., 2023) | - | 76.5 | 75.2 |
-| GeminiProVision+BDoG | - | 81.1 | 81.3 |
+| GPT-4+CoT (Wei et al., 2022) | 175B+ | 71.5 | 75.1 |
+| GPT-4+BDoG | 175B+ | 77.2 | 79.2 |
+| GeminiProVision (Team et al., 2023) | 175B+ | 76.5 | 75.2 |
+| GeminiProVision+BDoG | 175B+ | 81.1 | 81.3 |
+| Model | Method | ScienceQA-IMG-Dev | ScienceQA-IMG-Test | MMBench-Dev | ||||||||||||
+| NAT | SOC | LAN | Avg | NAT | SOC | LAN | Avg | LR | AR | RR | FP-S | FP-C | CP | Avg | ||
+| MniGPT-4 (Zhu et al., 2024) | Base | 42.9 | 30.6 | 43.7 | 38.4 | 42.0 | 30.1 | 50.0 | 37.7 | 7.5 | 31.3 | 4.3 | 30.3 | 9.0 | 35.6 | 24.3 |
+| Qwen-VL (Bai et al., 2023) | 52.1 | 59.8 | 58.3 | 55.0 | 55.7 | 62.0 | 77.3 | 58.7 | 16.1 | 44.7 | 34.8 | 35.2 | 39.2 | 46.6 | 38.2 | |
+| Qwen-VL-Chat (Bai et al., 2023) | 60.9 | 67.4 | 62.5 | 63.3 | 67.7 | 69.6 | 75.0 | 68.6 | 32.2 | 59.8 | 43.5 | 66.2 | 48.3 | 79.4 | 60.6 | |
+| mPLUG-Owl2 (Xu et al., 2023) | 60.6 | 68.0 | 45.8 | 62.8 | 62.5 | 66.2 | 61.4 | 63.9 | 32.2 | 72.4 | 60.9 | 68.6 | 60.1 | 79.4 | 66.5 | |
+| CogVLM-Chat (Wang et al., 2023c) | 63.1 | 69.2 | 77.1 | 65.6 | 68.0 | 72.2 | 70.4 | 69.7 | 29.7 | 65.8 | 60 | 66.9 | 58 | 76.7 | 63.7 | |
+| LLaVA-v1.5 (Liu et al., 2024) | 66.1 | 74.9 | 72.9 | 69.4 | 70.1 | 74.2 | 81.8 | 71.9 | 44.1 | 67.3 | 60.0 | 72.0 | 59.4 | 82.1 | 68.2 | |
+| InstructBLIP (Dai et al., 2023) | Base | 53.7 | 57.3 | 47.9 | 54.8 | 58.1 | 61.0 | 61.4 | 59.2 | 19.1 | 54.2 | 34.8 | 47.8 | 24.8 | 56.4 | 44.0 |
+| + BDoGDebate | 59.7 | 55.6 | 54.2 | 58.1 | 63.1 | 58.2 | 72.7 | 61.4 | 22.9 | 60.3 | 52.2 | 54.3 | 28.0 | 68.9 | 52.4 | |
+| + BDoGGraph | 58.1 | 61.3 | 52.1 | 59.0 | 60.6 | 62.6 | 68.2 | 61.5 | 58.8 | 65.5 | 41.2 | 51.2 | 18.6 | 46.1 | 51.1 | |
+| + BDoG | 61.1 | 64.0 | 52.1 | 61.9 | 61.1 | 66.5 | 75.0 | 63.5 | 63.3 | 71.9 | 37.8 | 56.3 | 20.3 | 59.1 | 55.8 | |
+| GeminiProVision (Team et al., 2023) | Base | 68.9 | 81.6 | 75.0 | 73.7 | 72.9 | 81.5 | 88.6 | 76.5 | 55.9 | 80.4 | 73.9 | 79.5 | 61.5 | 82.1 | 75.2 |
+| + BDoGDebate | 73.3 | 81.1 | 77.1 | 76.2 | 75.3 | 82.8 | 93.2 | 78.5 | 71.1 | 85.1 | 83.1 | 78.9 | 71.9 | 81.3 | 79.3 | |
+| + BDoGGraph | 69.8 | 84.8 | 87.5 | 75.6 | 74.7 | 86.8 | 88.6 | 79.6 | 75.0 | 84.5 | 80.7 | 81.4 | 73.0 | 83.6 | 80.7 | |
+| + BDoG | 73.6 | 86.2 | 85.4 | 78.4 | 76.6 | 87.4 | 93.2 | 81.1 | 74.0 | 84.8 | 83.4 | 81.3 | 73.7 | 84.4 | 81.3 | |
 5.3. Performance Comparison to SOTA Methods
-We evaluate the proposed method by by comparing it against two sets of state-of-the-art approaches as follows:
-Open-Source Multimodal LLMs with Relatively Moderate Parameters including MiniGPT-4 (Zhu et al., 2024), Qwen-VL and Qwen-VL- Chat (Bai et al., 2023), CogVLM-Chat (Wang et al., 2023c), mPLUG-Owl2 (Xu et al., 2023), LLaVA-v1.5 (Liu et al., 2024), and InstructBLIP (Dai et al., 2023).
-These models have parameter scales ranging from 7B to 17B
+In contrast to the few-shot methodology, which exhibits susceptibility to the specific examples selected for training, we have opted for the zero-shot setting.
+This approach circumvents potential biases introduced by a limited sample size, ensuring a more robust and generalizable model.
+We evaluate the proposed method by by comparing it against two sets of SOTA approaches as follows:
+Open-Source Multimodal LLMs with Relatively Moderate Parameters including MiniGPT-4 (Zhu et al., 2024), Qwen-VL and Qwen-VL- Chat (Bai et al., 2023), CogVLM-Chat (Wang et al., 2023c), mPLUG-Owl2 (Xu et al., 2023), LLaVA-v1.5 (Liu et al., 2024), and InstructBLIP (Dai et al., 2023), with parameter scales ranging from 7B to 17B.
 Closed-Source Multimodal LLMs with Large-Scale Parameters: GPT-3.5 (Wei et al., 2022), GPT-4V (Achiam et al., 2023) and GeminiProVision (Team et al., 2023).
 Following the general standard, GPT-3.5 and GPT-4 have been incorporated with the CoT (Wei et al., 2022) or DDCoT (Zheng et al., 2023b) (built based on image captioning results).
-These models are known for their parameter scales above 175B and are considered to have the best performance in most of the literature.
+These models are known for their parameter scales above 175B and are considered to have the best performance.
 The results are shown in Table 1.
@@
 This can be seen in the improvement achieved by InstructBLIP, which has experienced a boost of 4.3% and achieves an accuracy of 63.5% on SQA-IMG, comparable to that of GPT-3.5.
+Similar results can be found in LLaVA-v1.5 with BDoG which gains the 71.1% accuracy in MMBench, comparable to the GPT-4 model.
 BDoG reinforces the multimodal reasoning.
@@
 Even the open-source VL models of the former group achieves comparable performance to those of the latter one, with much smaller parameter scales.
-With BDoG, which reinforces multimodal reasoning by graph regulation, the performance of direct multimodal reasoning of InstructBLIP and GeminiProVision have been improved by 6.1% and 19.8% on the MMBench dataset.
-| Model | Method | ScienceQA-IMG-Dev | ScienceQA-IMG-Test | MMBench-Dev | ||||||||||||
-| NAT | SOC | LAN | Avg | NAT | SOC | LAN | Avg | LR | AR | RR | FP-S | FP-C | CP | Avg | ||
-| MniGPT-4 (Zhu et al., 2024) | Base | 42.9 | 30.6 | 43.7 | 38.4 | 42.0 | 30.1 | 50.0 | 37.7 | 7.5 | 31.3 | 4.3 | 30.3 | 9.0 | 35.6 | 24.3 |
-| Qwen-VL (Bai et al., 2023) | 52.1 | 59.8 | 58.3 | 55.0 | 55.7 | 62.0 | 77.3 | 58.7 | 16.1 | 44.7 | 34.8 | 35.2 | 39.2 | 46.6 | 38.2 | |
-| Qwen-VL-Chat (Bai et al., 2023) | 60.9 | 67.4 | 62.5 | 63.3 | 67.7 | 69.6 | 75.0 | 68.6 | 32.2 | 59.8 | 43.5 | 66.2 | 48.3 | 79.4 | 60.6 | |
-| mPLUG-Owl2 (Xu et al., 2023) | 60.6 | 68.0 | 45.8 | 62.8 | 62.5 | 66.2 | 61.4 | 63.9 | 32.2 | 72.4 | 60.9 | 68.6 | 60.1 | 79.4 | 66.5 | |
-| CogVLM-Chat (Wang et al., 2023c) | 63.1 | 69.2 | 77.1 | 65.6 | 68.0 | 72.2 | 70.4 | 69.7 | 29.7 | 65.8 | 60 | 66.9 | 58 | 76.7 | 63.7 | |
-| LLaVA-v1.5 (Liu et al., 2024) | 66.1 | 74.9 | 72.9 | 69.4 | 70.1 | 74.2 | 81.8 | 71.9 | 44.1 | 67.3 | 60.0 | 72.0 | 59.4 | 82.1 | 68.2 | |
-| InstructBLIP (Dai et al., 2023) | Base | 53.7 | 57.3 | 47.9 | 54.8 | 58.1 | 61.0 | 61.4 | 59.2 | 14.2 | 46.3 | 22.6 | 37.0 | 21.4 | 49.0 | 36.0 |
-| + BDoG | 59.7 | 55.6 | 54.2 | 58.1 | 63.1 | 58.2 | 72.7 | 61.4 | 43.7 | 57.8 | 31.5 | 42.7 | 19.5 | 39.1 | 42.6 | |
-| + BDoG | 58.1 | 61.3 | 52.1 | 59.0 | 60.6 | 62.6 | 68.2 | 61.5 | 58.8 | 65.5 | 41.2 | 51.2 | 18.6 | 46.1 | 51.1 | |
-| + BDoG | 61.1 | 64.0 | 52.1 | 61.9 | 61.1 | 66.5 | 75.0 | 63.5 | 63.3 | 71.9 | 37.8 | 56.3 | 20.3 | 59.1 | 55.8 | |
-| GeminiProVision (Team et al., 2023) | Base | 68.9 | 81.6 | 75.0 | 73.7 | 72.9 | 81.5 | 88.6 | 76.5 | 55.9 | 80.4 | 73.9 | 79.5 | 61.5 | 82.1 | 75.2 |
-| + BDoG | 73.3 | 81.1 | 77.1 | 76.2 | 75.3 | 82.8 | 93.2 | 78.5 | 71.1 | 85.1 | 83.1 | 78.9 | 71.9 | 81.3 | 79.3 | |
-| + BDoG | 69.8 | 84.8 | 87.5 | 75.6 | 74.7 | 86.8 | 88.6 | 79.6 | 75.0 | 84.5 | 80.7 | 81.4 | 73.0 | 83.6 | 80.7 | |
-| + BDoG | 73.6 | 86.2 | 85.4 | 78.4 | 76.6 | 87.4 | 93.2 | 81.1 | 74.0 | 84.8 | 83.4 | 81.3 | 73.7 | 84.4 | 81.3 | |
+With BDoG, which reinforces multimodal reasoning by graph regulation, the performance of direct multimodal reasoning of InstructBLIP and GeminiProVision have been improved by 6.1% and 11.8% on the MMBench dataset.
 5.4. Ablation Study
-In order to gain a comprehensive understanding of BDoG, we conduct an ablation study where we decompose BDoG into two variants:
-BDoG: we remove the graph regulation and constraints, resulting in a debate-only approach (i.e., vanilla multi-agent debate) for investigating the specific contribution of the debating component of BDoG.
-BDoG: we remove the debating rounds, resulting in a graph-based reasoning method for investigating the specific contribution of the graph regulation component of BDoG.
+In order to gain a comprehensive understanding of BDoG, we conduct an ablation study by decomposing BDoG into two variants:
+BDoGDebate: we remove the graph regulation and constraints, resulting in a debate-only approach (i.e., vanilla multi-agent debate) for investigating the specific contribution of the debating component of BDoG.
+BDoGGraph: we remove the debating rounds, resulting in a graph-based reasoning method for investigating the specific contribution of the graph regulation component of BDoG.
 Moreover, we analyze the performance of the two variants on the benchmarks by breaking it down into subcategories.
@@
 However, when considering specific categories, distinctions in the contributions of the debate and graph components become apparent.
-Impact of the debate component: BDoG demonstrates consistent improvements across both benchmarks with a debate-only setting, which encourages LLM agents to collaboratively refine and correct prior responses.
-For science questions, BDoG facilitates the model’s focus on specific errors, such as direction, size, and position, leading to improved performance in the natural science domain (boosting accuracy from 53.7 to 59.7 for InstructBLIP and 68.9 to 73.3 for GeminiProVision).
-Additionally, BDoG significantly enhances logical reasoning tasks (improving accuracy from 14.2 to 43.7).
+Impact of the debate component: BDoGDebate demonstrates consistent improvements across both benchmarks with a debate-only setting, which encourages LLM agents to collaboratively refine and correct prior responses.
+For science questions, BDoGDebate facilitates the model’s focus on specific errors, such as direction, size, and position, leading to improved performance in the natural science domain (boosting accuracy from 53.7 to 59.7 for InstructBLIP and 68.9 to 73.3 for GeminiProVision).
 However, the debate-only nature has limitations, including trivialization and focus diversion issues.
-Without the graph regulation, overall performance decreases from 55.8 to 42.6 for InstructBLIP, particularly when addressing questions that require attention to specific attributes and relations.
-Impact of the graph regulation: With a graph-regularized knowledge base for the discussion, BDoG also demonstrates consistent improvement of overs the base models on both benchmarks.
-Compared to the text-based and debate-only method BDoG, it performs evidently better on the MMBench by .
-This is not a surprise, because MMBench emphasizes more the attribute and relations which challenges a model’s ability on structural reasoning.
-Although incorporating fact-related graph information proves beneficial in BDoG, the absence of the iteratively refined debate procedure results in decreased performance due to the coarse and distorted extraction of blueprint information.
+Without the graph regulation, overall performance decreases from 55.8 to 52.4 for InstructBLIP, particularly when addressing questions that require attention to multi-hop logistic reasoning (LR) and specific attributes (AR).
+Impact of the graph regulation: With a graph-regularized knowledge base for the discussion, BDoGGraph also demonstrates consistent improvement of overs the base models on both benchmarks.
+Compared to the text-based and debate-only method BDoGDebate, it performs evidently better on the logistic reasoning and attributes reasoning questions by addressing the opinion trivialization and diversion with initialized blueprint.
+Although incorporating fact-related graph information proves beneficial in BDoGGraph, the absence of the iteratively refined debate procedure results in decreased performance due to the coarse and distorted extraction of blueprint information.
 Impact of combining the debate and graph components: By combining the two components, BDoG achieves gains across nearly all categories.
@@
 This suggests that BDoG is robust and generalizes well for science-related questions.
-Remarkably, BDoG significantly outperforms the baseline model (InstructBLIP) on the MMBench-Dev set, particularly in the areas of Logical Reasoning (LR) with a margin of 49.1%, Attribute Reasoning (AR) with a margin of 26.6%, and Relation Reasoning (RR) with a margin of 15.2%.
+Remarkably, BDoG significantly outperforms the baseline model (InstructBLIP) on the MMBench-Dev set, particularly in the areas of Logical Reasoning (LR) with a margin of 44.2%, Attribute Reasoning (AR) with a margin of 17.7%, and Relation Reasoning (RR) with a margin of 3%.
 BDoG enhances logical reasoning (LR) through a mechanism that refines the reasoning process iteratively, emphasizing the importance of multi-step reasoning rationales.
@@
 This is evident in the consistent improvement observed on the blueprint graph, showcasing the combined benefits of these two components.
-Figure 3 provides running examples demonstrating the superior reasoning performance of our proposed BDoG framework compared to the BDoG method.
+Figure 3 provides running examples demonstrating the superior reasoning performance of our proposed BDoG framework compared to the BDoGDebate method.
 The left case draws from the ScienceQA dataset, testing geographic knowledge and map interpretation.
-While BDoG correctly answered Dominica is highlighted, it also generated irrelevant information about Dominica’s economic development.
+While BDoGDebate correctly answered Dominica is highlighted, it also generated irrelevant information about Dominica’s economic development.
 This misguided the agents into off-topic discussion, concluding incorrectly with Barbados.
@@
 As the image contained both candies and jars, it posed a challenge.
-With BDoG relying on text alone, agreement was rarely reached as responses changed over debate rounds.
+With BDoGDebate relying on text alone, agreement was rarely reached as responses changed over debate rounds.
 However, BDoG first generated a blueprint defining image objects and attributes.
@@
 It output the final answer by comparing and connecting the two jar sub-graphs.
-In summary, Figure 3 demonstrates that BDoG beats BDoG on both datasets through its blueprint-driven approach.
-This concentrates graph-based reasoning on salient topics and prunes irrelevant details to arrive at well-supported conclusions.
 | Round | ScienceQA-IMG-Test | MMBench-Dev | ||
@@
 5.5. Monitoring The Debating Progress
-We evaluated the model’s performance against the termination criteria across multiple debate rounds based on the data in Table 3.
-Our analysis showed that for models with smaller parameters like InstructBLIP, moving from a single round to two rounds led to significant gains in performance.
-This improvement was particularly notable when increasing the number of rounds from one to two.
-However, for larger models that may reach agreement more easily, the performance enhancement was relatively modest when amplifying the number of debate rounds.
-In general, we found the model’s performance tended to converge within the second or third round.
+We evaluate the model’s performance against the termination criteria across multiple debate rounds based on the data in Table 3.
+Our analysis shows that for models with smaller parameters like InstructBLIP, moving from a single round to two rounds led to significant gains in performance.
+This improvement is particularly notable when increasing the number of rounds from one to two.
+However, for larger models that may reach agreement more easily, the performance enhancement is relatively modest when amplifying the number of debate rounds.
+In general, we find the model’s performance tend to converge within the second or third round.
 This can be attributed to the underlying reasoning typically being able to answer questions within 2-3 steps.
@@
 5.6. Efficiency Analysis
-We further compared the effectiveness versus efficiency of our BDoG framework against BDoG, with results shown in Figure 5.
-Maintaining concise content focused on key aspects, the graph structure of BDoG demonstrated superior efficiency, requiring approximately 50% less inference time than BDoG.
+We further compare the effectiveness versus efficiency of our BDoG framework against BDoGDebate, as shown in Figure 5.
+Maintaining concise content focuses on key aspects, the graph structure of BDoG demonstrates superior efficiency, requiring approximately 50% less inference time than BDoGDebate.
 By first generating a blueprint, BDoG defines the scope of the current state, thereby improving model efficiency by filtering irrelevant information.
-Concurrently, Figure 5 shows BDoG outperformed BDoG in effectiveness, achieving over 5 percentage points higher accuracy than BDoG across three test sets.
-This enhanced effectiveness can be attributed to BDoG’s framework concentrating on important knowledge rather than generational textual content without guidance, as in BDoG.
-Defining a blueprint to establish the discussion boundaries before generating a graph structure enhances both efficiency and effectiveness by restricting the model’s reasoning to salient topics.
-In summary, Figure 5 illustrates BDoG demonstrated better time performance and predictive power compared to BDoG, validating the utility of our approach in balancing computational cost and predictive accuracy through focused, blueprint-driven debate.
+Concurrently, Figure 5 shows BDoG outperforms BDoGDebate in effectiveness, achieving over 5 percentage higher accuracy than BDoGDebate across three test sets.
+This enhanced effectiveness can be attributed to BDoG’s concentrating on salient knowledge rather than generational textual content without guidance, as in BDoGDebate.
 6. Conclusion
 This paper has presented a pioneering pilot study that introduces multi-agent debate into the realm of multimodal reasoning.
-We tackled two prominent challenges faced in this context: the issue of opinions being trivialized and focus diversion.
-By recognizing the limitations of existing debating schemes, we propose Blueprint Debate on Graphs (BDoG), which confines debates to a blueprint graph and stores evidence in graph branches, to address the challenges of word-level opinion trivialization and distraction caused by irrelevant concepts.
+We propose Blueprint Debate on Graphs (BDoG), which confines debates to a blueprint graph and stores evidence in graph branches, to address the challenges of word-level opinion trivialization and distraction caused by irrelevant concepts.
 Extensive experiments conducted in ScienceQA and MMBench validate the efficacy of BDoG, surpassing previous methods and establishing new state-of-the-art results.
+Acknowledgements.
+This research was supported by the HK RGC Theme-based Research Scheme (PolyU No.: T43-513/23-N) and the National Natural Science Foundation of China (Grant No.: 62372314).References
+- (1)
+- Achiam et al.
+(2023) Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, Shyamal Anadkat, et al.
+2023.
+Gpt-4 technical report.
+arXiv preprint arXiv:2303.08774 (2023).
+- Bai et al.
+(2023) Jinze Bai, Shuai Bai, Shusheng Yang, Shijie Wang, Sinan Tan, Peng Wang, Junyang Lin, Chang Zhou, and Jingren Zhou.
+2023.
+Qwen-vl: A frontier large vision-language model with versatile abilities.
+arXiv preprint arXiv:2308.12966 (2023).
+- Chan et al.
+(2023) Chi-Min Chan, Weize Chen, Yusheng Su, Jianxuan Yu, Wei Xue, Shanghang Zhang, Jie Fu, and Zhiyuan Liu.
+2023.
+ChatEval: Towards Better LLM-based Evaluators through Multi-Agent Debate.
+In The Twelfth International Conference on Learning Representations.
+- Chen et al.
+(2015) Xinlei Chen, Hao Fang, Tsung-Yi Lin, Ramakrishna Vedantam, Saurabh Gupta, Piotr Dollár, and C Lawrence Zitnick.
+2015.
+Microsoft coco captions: Data collection and evaluation server.
+arXiv preprint arXiv:1504.00325 (2015).
+- Cobbe et al.
+(2021) Karl Cobbe, Vineet Kosaraju, Mohammad Bavarian, Mark Chen, Heewoo Jun, Lukasz Kaiser, Matthias Plappert, Jerry Tworek, Jacob Hilton, Reiichiro Nakano, et al.
+2021.
+Training verifiers to solve math word problems.
+arXiv preprint arXiv:2110.14168 (2021).
+- Dai et al.
+(2023) Wenliang Dai, Junnan Li, Dongxu Li, Anthony Meng Huat Tiong, Junqi Zhao, Weisheng Wang, Boyang Albert Li, Pascale Fung, and Steven C.
+H.
+Hoi.
+2023.
+InstructBLIP: Towards General-purpose Vision-Language Models with Instruction Tuning.
+ArXiv abs/2305.06500 (2023).
+- Du et al.
+(2023) Yilun Du, Shuang Li, Antonio Torralba, Joshua B Tenenbaum, and Igor Mordatch.
+2023.
+Improving Factuality and Reasoning in Language Models through Multiagent Debate.
+arXiv preprint arXiv:2305.14325 (2023).
+- Goyal et al.
+(2017) Yash Goyal, Tejas Khot, Douglas Summers-Stay, Dhruv Batra, and Devi Parikh.
+2017.
+Making the v in vqa matter: Elevating the role of image understanding in visual question answering.
+In Proceedings of the IEEE conference on computer vision and pattern recognition.
+6904–6913.
+- He et al.
+(2020) Jie He, Tao Wang, Deyi Xiong, and Qun Liu.
+2020.
+The box is in the pen: Evaluating commonsense reasoning in neural machine translation.
+In Findings of the Association for Computational Linguistics: EMNLP 2020.
+3662–3672.
+- Hendrycks et al.
+(2020) Dan Hendrycks, Collin Burns, Steven Basart, Andy Zou, Mantas Mazeika, Dawn Song, and Jacob Steinhardt.
+2020.
+Measuring Massive Multitask Language Understanding.
+In International Conference on Learning Representations.
+- Horawalavithana et al.
+(2023) Sameera Horawalavithana, Sai Munikoti, Ian Stewart, and Henry Kvinge.
+2023.
+Scitune: Aligning large language models with scientific multimodal instructions.
+arXiv preprint arXiv:2307.01139 (2023).
+- Hu et al.
+(2023) Linmei Hu, Zeyi Liu, Ziwang Zhao, Lei Hou, Liqiang Nie, and Juanzi Li.
+2023.
+A survey of knowledge enhanced pre-trained language models.
+IEEE Transactions on Knowledge and Data Engineering (2023).
+- Huang et al.
+(2023) Jie Huang, Xinyun Chen, Swaroop Mishra, Huaixiu Steven Zheng, Adams Wei Yu, Xinying Song, and Denny Zhou.
+2023.
+Large Language Models Cannot Self-Correct Reasoning Yet.
+In The Twelfth International Conference on Learning Representations.
+- Li et al.
+(2023b) Junnan Li, Dongxu Li, Silvio Savarese, and Steven Hoi.
+2023b.
+Blip-2: Bootstrapping language-image pre-training with frozen image encoders and large language models.
+arXiv preprint arXiv:2301.12597 (2023).
+- Li et al.
+(2023a) Kun Li, Jiaxiu Li, Dan Guo, Xun Yang, and Meng Wang.
+2023a.
+Transformer-based visual grounding with cross-modality interaction.
+ACM Transactions on Multimedia Computing, Communications and Applications 19, 6 (2023), 1–19.
+- Li et al.
+(2023c) Xingxuan Li, Ruochen Zhao, Yew Ken Chia, Bosheng Ding, Shafiq Joty, Soujanya Poria, and Lidong Bing.
+2023c.
+Chain-of-Knowledge: Grounding Large Language Models via Dynamic Knowledge Adapting over Heterogeneous Sources.
+In The Twelfth International Conference on Learning Representations.
+- Liang et al.
+(2023) Tian Liang, Zhiwei He, Wenxiang Jiao, Xing Wang, Yan Wang, Rui Wang, Yujiu Yang, Zhaopeng Tu, and Shuming Shi.
+2023.
+Encouraging Divergent Thinking in Large Language Models through Multi-Agent Debate.
+arXiv preprint arXiv:2305.19118 (2023).
+- Liu et al. (2024) Haotian Liu, Chunyuan Li, Qingyang Wu, and Yong Jae Lee. 2024. Visual instruction tuning. Advances in neural information processing systems 36 (2024).
+- Liu et al.
+(2023) Yuan Liu, Haodong Duan, Yuanhan Zhang, Bo Li, Songyang Zhang, Wangbo Zhao, Yike Yuan, Jiaqi Wang, Conghui He, Ziwei Liu, et al.
+2023.
+Mmbench: Is your multi-modal model an all-around player?
+arXiv preprint arXiv:2307.06281 (2023).
+- Lu et al.
+(2022) Pan Lu, Swaroop Mishra, Tanglin Xia, Liang Qiu, Kai-Wei Chang, Song-Chun Zhu, Oyvind Tafjord, Peter Clark, and Ashwin Kalyan.
+2022.
+Learn to explain: Multimodal reasoning via thought chains for science question answering.
+Advances in Neural Information Processing Systems 35 (2022), 2507–2521.
+- Lu et al.
+(2024) Pan Lu, Baolin Peng, Hao Cheng, Michel Galley, Kai-Wei Chang, Ying Nian Wu, Song-Chun Zhu, and Jianfeng Gao.
+2024.
+Chameleon: Plug-and-play compositional reasoning with large language models.
+Advances in Neural Information Processing Systems 36 (2024).
+- Madaan et al.
+(2024) Aman Madaan, Niket Tandon, Prakhar Gupta, Skyler Hallinan, Luyu Gao, Sarah Wiegreffe, Uri Alon, Nouha Dziri, Shrimai Prabhumoye, Yiming Yang, et al.
+2024.
+Self-refine: Iterative refinement with self-feedback.
+Advances in Neural Information Processing Systems 36 (2024).
+- Mitra et al.
+(2023) Chancharik Mitra, Brandon Huang, Trevor Darrell, and Roei Herzig.
+2023.
+Compositional chain-of-thought prompting for large multimodal models.
+arXiv preprint arXiv:2311.17076 (2023).
+- Mondal et al.
+(2024) Debjyoti Mondal, Suraj Modi, Subhadarshi Panda, Rituraj Singh, and Godawari Sudhakar Rao.
+2024.
+KAM-CoT: Knowledge Augmented Multimodal Chain-of-Thoughts Reasoning.
+arXiv preprint arXiv:2401.12863 (2024).
+- Shinn et al.
+(2024) Noah Shinn, Federico Cassano, Ashwin Gopinath, Karthik Narasimhan, and Shunyu Yao.
+2024.
+Reflexion: Language agents with verbal reinforcement learning.
+Advances in Neural Information Processing Systems 36 (2024).
+- Sun et al.
+(2023) Jiashuo Sun, Chengjin Xu, Lumingyuan Tang, Saizhuo Wang, Chen Lin, Yeyun Gong, Lionel Ni, Heung-Yeung Shum, and Jian Guo.
+2023.
+Think-on-Graph: Deep and Responsible Reasoning of Large Language Model on Knowledge Graph.
+In The Twelfth International Conference on Learning Representations.
+- Team et al.
+(2023) Gemini Team, Rohan Anil, Sebastian Borgeaud, Yonghui Wu, Jean-Baptiste Alayrac, Jiahui Yu, Radu Soricut, Johan Schalkwyk, Andrew M Dai, Anja Hauth, et al.
+2023.
+Gemini: a family of highly capable multimodal models.
+arXiv preprint arXiv:2312.11805 (2023).
+- Wang et al.
+(2023a) Keheng Wang, Feiyu Duan, Sirui Wang, Peiguang Li, Yunsen Xian, Chuantao Yin, Wenge Rong, and Zhang Xiong.
+2023a.
+Knowledge-driven cot: Exploring faithful reasoning in llms for knowledge-intensive question answering.
+arXiv preprint arXiv:2308.13259 (2023).
+- Wang et al.
+(2023b) Lei Wang, Yi Hu, Jiabang He, Xing Xu, Ning Liu, Hui Liu, and Heng Tao Shen.
+2023b.
+T-SciQ: Teaching Multimodal Chain-of-Thought Reasoning via Large Language Model Signals for Science Question Answering.
+arXiv preprint arXiv:2305.03453 (2023).
+- Wang et al.
+(2023c) Weihan Wang, Qingsong Lv, Wenmeng Yu, Wenyi Hong, Ji Qi, Yan Wang, Junhui Ji, Zhuoyi Yang, Lei Zhao, Xixuan Song, et al.
+2023c.
+Cogvlm: Visual expert for pretrained language models.
+arXiv preprint arXiv:2311.03079 (2023).
+- Wang et al.
+(2023d) Yanan Wang, Michihiro Yasunaga, Hongyu Ren, Shinya Wada, and Jure Leskovec.
+2023d.
+VQA-GNN: Reasoning with Multimodal Knowledge via Graph Neural Networks for Visual Question Answering.
+In Proceedings of the IEEE/CVF International Conference on Computer Vision.
+21582–21592.
+- Wei et al.
+(2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al.
+2022.
+Chain-of-thought prompting elicits reasoning in large language models.
+Advances in Neural Information Processing Systems 35 (2022), 24824–24837.
+- Welleck et al.
+(2022) Sean Welleck, Ximing Lu, Peter West, Faeze Brahman, Tianxiao Shen, Daniel Khashabi, and Yejin Choi.
+2022.
+Generating Sequences by Learning to Self-Correct.
+In The Eleventh International Conference on Learning Representations.
+- Wu et al.
+(2020) Zhiwei Wu, Changmeng Zheng, Yi Cai, Junying Chen, Ho-fung Leung, and Qing Li.
+2020.
+Multimodal representation with embedded visual guiding objects for named entity recognition in social media posts.
+In Proceedings of the 28th ACM International conference on multimedia.
+1038–1046.
+- Xu et al.
+(2023) Haiyang Xu, Qinghao Ye, Mingshi Yan, Yaya Shi, Jiabo Ye, Yuanhong Xu, Chenliang Li, Bin Bi, Qiuchen Qian, Wei Wang, Guohai Xu, Ji Zhang, Songfang Huang, Feiran Huang, and Jingren Zhou.
+2023.
+mPLUG-2: A Modularized Multi-modal Foundation Model Across Text, Image and Video.
+In International Conference on Machine Learning.
+- Yin et al.
+(2023) Zhangyue Yin, Qiushi Sun, Cheng Chang, Qipeng Guo, Junqi Dai, Xuan-Jing Huang, and Xipeng Qiu.
+2023.
+Exchange-of-thought: Enhancing large language model capabilities through cross-model communication.
+In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing.
+15135–15153.
+- Zhang et al. (2023a) Jintian Zhang, Xin Xu, and Shumin Deng. 2023a. Exploring collaboration mechanisms for llm agents: A social psychology view. arXiv preprint arXiv:2310.02124 (2023).
+- Zhang et al.
+(2023b) Zhuosheng Zhang, Aston Zhang, Mu Li, Hai Zhao, George Karypis, and Alex Smola.
+2023b.
+Multimodal chain-of-thought reasoning in language models.
+arXiv preprint arXiv:2302.00923 (2023).
+- Zhao and Wu (2023) Wentian Zhao and Xinxiao Wu. 2023. Boosting entity-aware image captioning with multi-modal knowledge graph. IEEE Transactions on Multimedia (2023).
+- Zheng et al.
+(2021) Changmeng Zheng, Junhao Feng, Ze Fu, Yi Cai, Qing Li, and Tao Wang.
+2021.
+Multimodal relation extraction with efficient graph alignment.
+In Proceedings of the 29th ACM international conference on multimedia.
+5298–5306.
+- Zheng et al.
+(2023a) Chuanyang Zheng, Zhengying Liu, Enze Xie, Zhenguo Li, and Yu Li.
+2023a.
+Progressive-hint prompting improves reasoning in large language models.
+arXiv preprint arXiv:2304.09797 (2023).
+- Zheng et al.
+(2023b) Ge Zheng, Bin Yang, Jiajin Tang, Hong-Yu Zhou, and Sibei Yang.
+2023b.
+DDCoT: Duty-Distinct Chain-of-Thought Prompting for Multimodal Reasoning in Language Models.
+In Thirty-seventh Conference on Neural Information Processing Systems.
+- Zhu et al.
+(2024) Deyao Zhu, Jun Chen, Xiaoqian Shen, Xiang Li, and Mohamed Elhoseiny.
+2024.
+MiniGPT-4: Enhancing Vision-Language Understanding with Advanced Large Language Models.
+In The Twelfth International Conference on Learning Representations.
+https://openreview.net/forum?id=1tZbq88f27
+Appendix A Experimental Details
+A.1. Algorithm for BDoG
+For a better understanding of BDoG, an algorithmic procedure has been formulated in Algorithm 1.
+As depicted in Figure 6, we introduce the Blueprint debate-on-graph (BDoG) paradigm modeled by a hierarchical tree topology.
+The topology is defined by the vertex set containing nodes and edge set representing connections.
+Within the BDoG framework, leaf nodes directly exchange proposed updates to the shared knowledge graph, while interior nodes aggregate information flows.
+Graph condensation aims to learn a small, synthetic, and informative graph from a large, original graph .
+We consider the graph condensation as a reasoning process that multimodal LLMs learn to revise previous errors and filter out irrelevant information.
+We formally define the condensation process as follows:
+Entity update: Let and represent the sets of entities in graphs and respectively after a debate round.
+The set of common agreed-upon entities is defined as:
+Relation update: Let and represent the sets of relations in graphs and .
+The set of common agreed-upon relations is defined as:
+Graph pruning: After identifying the common agreed-upon entities () and relations (), the LLM is prompted to discard any entities and relations in graphs and that are deemed irrelevant to the discussion.
+The LLM evaluates each entity and relation not in the agreed set, and determines whether to keep it based on its relevance to the problem context and debate.
+Any deemed irrelevant are removed.
+The same process occurs for pruning .
+This prompts the LLM to actively discard unnecessary information based on learned relevance, rather than a rigid mathematical operation, better simulating the flexibility of human judgment.
+A.2. Statistics of Datasets
+Table 4 provides an overview of the size and diversity of datasets used in the paper, including the number of instances, subjects, categories, and the average question length.
+These statistics can help in understanding the complexity and challenges posed by these datasets for multimodal reasoning-based QA systems.
+| Dataset | Instance | Subject | Category | Avg. Ques. |
+| SQA-Test (Lu et al., 2022) | 2017 | 3 | 65 | 9.3 |
+| SQA-Dev (Lu et al., 2022) | 2097 | 3 | 66 | 9.6 |
+| MMBench-Dev (Liu et al., 2023) | 4329 | 6 | 20 | 8.9 |
+| Setting | Value |
+| LLM | Vicuna-13B |
+| Vision Encoder | EVA CLIP-G/14 |
+| Hardware Requirement | 2x A100 (40GB) |
+| Truncation Mode | Left |
+| Number of Beams | 5 |
+| Temperature | 1.0 |
+| Top-p | 0.9 |
+| Data Type | float32 |
+| Image Resolution | 224x224 |
+| Maximum Input Length | 256 |
+| Maximum Output Graph Length | 128 |
+| Maximum Output Answer Length | 50 |
+| Maximum Debate Round | 4 |
+| Inference Time for SQA | 4.7 s/sample |
+| Inference Time for MMBench | 5.4 s/sample |
+A.3. Model Deployment
+The specifics of model deployment and hyperparameter configurations for the InstructBLIP system are detailed in Table 5.
+Experimental evaluations are conducted leveraging the computational resources of two NVIDIA A100 GPUs.
+Consistent with prior work (Dai et al., 2023), we employ the Vicuna-13B as the large language model and the EVA-CLIP as the vision encoding module.
+It is noteworthy that InstructBLIP imposes constraints on the overall input length; consequently, we set the output limits for graph generation and candidate answer production to 128 and 50 tokens, respectively.
+For outputs exceeding 256 tokens in length, we apply truncation techniques.
+Furthermore, we report the inference time metrics for the ScienceQA (SQA) and MMBench datasets.
+Although the average question token count for MMBench is lower than SQA, the inference time required is higher.
+A plausible explanation for this discrepancy may be the inherently greater complexity of questions in the MMBench dataset, necessitating the generation of more intricate output graphs.
+For the GeminiProVision and GPT-4V systems, we utilize their official APIs without employing any pre-processing or post-processing techniques.
+A.4. Prompts
+A.4.1. Role Specification
+To simulate a debate process, we treat each multimodal LLM as an agent and specify distinct roles to them following previous studies.
+Although all the agents share similar actions defined by several instructions, formulate an analogy personality for each agent can encourage exchange of thoughts from diverse background and knowledge sources.
+Inspired by this, we explicitly model different roles as follows:
+The diligent explainer:
+Role = Scrutinize (Problem) + Explain (Details, Knowledge, Logic)
+Goal = Benchmark (Solution)
+[hidealllines=true,backgroundcolor=gray!20]
+You are Ben, a high school student with a track record of excellent grades, particularly in mathematics.
+Your friends admire your diligence and often seek your guidance in their studies.
+Your role is to scrutinize the problem at hand with your usual attention to detail, drawing from your vast knowledge of principles.
+Your clear and logical explanations are valuable, as they will serve as a benchmark for your friends to compare and refine their own solutions.
+The creative problem-solver:
+Role = Dissect (Problem) + Leverage (Creative Strategies)
+Goal = Devise (Unique Solution)
+[hidealllines=true,backgroundcolor=gray!20]
+You are Peter, a high school student recognized for your unique problem-solving abilities.
+Your peers often turn to you for assistance when they encounter challenging tasks, as they appreciate your knack for devising creative solutions.
+Today, your challenge is to dissect the given problem, leveraging your unique problem-solving strategies.
+The attentive analyzer:
+Role = Analyze (Problem) + Apply (Attentive Skills) + PieceTogether (Detailed Solution)
+Goal = Catch (Missed Details)
+[hidealllines=true,backgroundcolor=gray!20]
+You are Kitty, a high school student admired for your attentiveness and detail-oriented nature.
+Your friends often rely on you to catch details they might have missed in their work.
+Your task is to carefully analyze the presented problem, apply your attentive skills, and piece together a detailed solution.
+In these expressions, Scrutinize, Dissect, Analyze represent the core activities of examining the problem, Details, Knowledge, Logic, Creative Strategies, Attentive Skills represent individual traits, and the ”+” signifies combining activities and goals.
+A.4.2. Blueprint Extraction
+Given an question and its correlated images , BDoG formulates a blueprint encompassing coarse-grained descriptive information, thereby circumventing the necessity for manually annotated ground-truth scene graphs or image captions.
+To relate the visual and textual contexts, we initially investigate disaggregating the graph extraction process into three dissociated components: (1) Generation of captions depicting .
+(2) Detection of semantic associations between and .
+(3) Aggregation of and to compose representing the unified multimodal representation, where nodes signify detected visual objects and edges capture inter-object relations as informed by along with image-question alignments from .
+Although this approach shows potential, such a decomposed pipeline is susceptible to the error propagation problem.
+The quality of the generated graph will be influenced by introduced biases and erroneous captions pertaining to specific objects.
+To mitigate this issue, we aim to adopt a holistic perspective not only of the objects, which serve as primary units for visual reasoning, but also of their properties and interactions in a “big picture” manner visualized simultaneously.
+Specifically, we prompt the multimodal large language model to generate a graph which consists of objects , attributes and relationships that are relevant to answering the question .
+[hidealllines=true,backgroundcolor=gray!20]
+Hint: {context} Question: {question} Options: {choices}
+For the provided image and its associated question. generate a scene graph in JSON format that includes the following:
+1. Objects that are relevant to answering the question.
+2. Object attributes that are relevant to answering the question.
+3. Object relationships that are relevant to answering the question.
+A.4.3. Proponent and Opponent
+The proponent and opponent are initialized by requiring them to adhere to the constraints for graph condensation, which involves updating, adding, and pruning nodes.
+The prompt for proponent is as the following:
+[hidealllines=true,backgroundcolor=gray!20]
+You are a fellow debater from the AFFIRMATIVE side.
+For the provided image and its associated question, generate an updated graph from a different view based on the Debate Graph in JSON format that includes the following:
+1. Objects that are more relevant to answering the question.
+2. Object attributes that are more relevant to answering the question.
+3. Object relationships that are more relevant to answering the question.
+4. Delete the irrelevant objects, attributes and relationships
+Hint: {context} Question: {question} Options: {choices} Debate Graph: {knowledge} Updated Graph:
+Similarly, the prompt for opponent can be implemented as:
+[hidealllines=true,backgroundcolor=gray!20]
+You are a fellow debater from the NEGATIVE side.
+For the provided image and its associated question, generate an updated graph from a different view based on the Debate Graph in JSON format that includes the following:
+1. Objects that are more relevant to answering the question.
+2. Object attributes that are more relevant to answering the question.
+3. Object relationships that are more relevant to answering the question.
+4. Delete the irrelevant objects, attributes and relationships
+Hint: {context} Question: {question} Options: {choices} Debate Graph: {knowledge} Updated Graph:
+It is important to note that the input debate graph for the proponent can be the updated graph from the opponent, and vice versa.
+A.4.4. Moderator
+We contend that prior research has employed an iterative approach involving text summarization to refine output, which has led to the trivialization of opinions characterized by a decline in performance.
+In contrast to this approach, our methodology directly extracts the answer through the analysis of the final graphs generated by both the proponent and opponent arguments.
+The prompt is presented as follows:
+[hidealllines=true,backgroundcolor=gray!20]
+Hint: {context} Affirmative Graph: {knowledge[0]} Negative Graph : {knowledge[1]}
+Use the image and two debate graph as context and answer the following question: Question: {question} Options: {choices}
+Answer with the option’s letter from the given choices directly.
+Notably, visual elements are seamlessly integrated throughout reasoning processes, ensuring that the textual output is grounded in the visual context.
+Appendix B Qualitative Examples
+Figure 7 depicts the running examples within a debate round that compares BDoG with MAD on ScienceQA-IMG and MMBench datasets.
+Our proposed Blueprint debate on graph (BDoG) is a more effective way to present information than the vanilla Multi-agent debate (MAD).
+It is more structured, more visual, and more interactive.
+This makes it easier to follow the flow of the debate, to identify the key points that are being made, and to explore the information in more detail.
+The first case utilizes the ScienceQA dataset, evaluating geographic knowledge and map interpretation skills.
+The correct response (identifying the southernmost state) necessitates comparing the locations of various states.
+In the multi-agent debate scenario, despite the affirmative agent’s accurate prediction of the relative position, the negative side overwhelmingly opposes it.
+Due to the widespread issue of hallucination in large language models (LLMs), detecting such misinformation demands significant effort.
+Instead of engaging in debate solely on the final answer, our model facilitates debate at the fact level.
+By strategically modifying blueprint nodes, BDoG demonstrates that the final answer should be derived by comparing the attributes (latitudes) of each candidate.
+BDoG offers the additional advantage of facilitating easier monitoring of changes, resulting in a more reliable system.
+In the second case, drawn from the MMBench dataset, the task requires reasoning about attributes to answer the posed question.
+While the MAD method generates rationales incorporating extensive inherent knowledge and imagination, this leads to the erroneous inference of a live Ursus, distinct from the fossil depicted in the image.
+Notably, the question demands direct observation of the image, from which the correct answer – ”Long legs” – can be readily inferred.
+The extraneous information generated by MAD misguides the model towards irrelevant concepts.
+Conversely, our BDoG method commences by analyzing the image, the associated question, and the candidate options.
+This effectively restricts the scope of analysis to the attributes of the fossil.
+Subsequently, BDoG incorporates additional observed features and refines potentially inaccurate nodes, ultimately leading to the accurate prediction.
+Appendix C Other Essentials of the Model
+C.1. Monitoring the Debate Progress
+Figures 8 and 9 depict the evolution of debate graphs for the ScienceQA-IMG dev set and MMBench-Dev set, respectively.
+These figures illustrate the dynamic changes in attributes, entities, and relations across debate rounds.
+Notably, the proposed BDoG framework outputs debate graphs in JSON format, facilitating further analysis.
+By comparing updated debate graphs with their original counterparts, we can track the modifications introduced during the debate process.
+This analysis underpins a key strength of BDoG: its ability to quantify the debate process through the lens of graph changes.
+This approach effectively demonstrates the value of dynamically adjusting the initial graph based on the evolving discussion.
+Furthermore, the results presented in Figures 8 and 9 align with our hypothesis that disagreements and errors diminish as the debate progresses.
+0pt
+olFFL
+0pt
+olFFL
+C.2. Effect of Blueprint Quality
+We conduct a human evaluation to assess the impact of blueprint quality, as illustrated in Figure 10.
+A random sample of 200 predictions from the MMBench dataset is selected for evaluation.
+Due to the absence of a standardized metric for evaluating generated graph quality, three annotators are tasked with classifying each blueprint as either high or low quality.
+The results reveal that 56% of the initial blueprints were classified as low-quality.
+This finding aligns with expectations, given the complexity of the questions and the potential for the MLLM to generate coarse and imprecise direct answers.
+To address the limitations of low-quality blueprints, we propose BDoG, a novel approach that iteratively refines the blueprint to enhance its conciseness and ultimately converge towards a correct answer.
+As demonstrated in the left panel of Figure 10, the final correctness rate is strongly correlated with blueprint quality.
+Notably, for questions with high-quality initial blueprints, the final correctness rate reaches 93.2%, highlighting the importance of a well-constrained initial graph.
+Furthermore, it is noteworthy that 67.8% of instances with low-quality blueprints ultimately result in correct predictions, demonstrating the effectiveness of BDoG’s iterative refinement capabilities.
+C.3. Effect of Introduced Options
+We further investigate the impact of incorporating options within a debate round.
+For a moderately parameterized multimodal LLM such as InstructBLIP-13B, introducing options can introduce noise and degrade final prediction performance.
+This is likely due to the model’s inherent under-confidence in its judgments, leading to options becoming a distraction that hinders effective reasoning.
+By ablating the options from the debate round, the performance of the InstructBLIP-13B model improves by 2.4% to 3.4%.
+However, this improvement is relatively minor for the larger GeminiProVision model.
+Such multimodal LLMs with over 175B parameters exhibit greater robustness to distractions from options, likely due to their increased capacity and sophistication.